# Identifying Reliable Predictions in Detection Transformers

### v1

@@ Appendix G Comparative Study on Post-Processing
Table 6: Comparison of the proposed method ($\mathsf{ContrastiveConf}$) with baseline methods on their Pearson correlation coefficient with the image-level reliability in Equation (6), with different models. $\mathsf{Perf}$ is evaluated using AP (top), AP50 (middle), and AP75 (bottom) metrics. For each model, we apply the optimal matching, the standard post-processing scheme (Top-$100$ and confidence thresholding by $0.3$), and the proposed OCE-based post-processing scheme, as indicated within the parentheses. The strongest correlations are highlighted in bold, while negative correlations appear in red. First, the proposed method achieves the highest correlation with image-level reliability. In addition, it is noteworthy that the average confidence of negative predictions ($\mathsf{Conf}^{-}$) exhibits a negative correlation with image-level reliability, in contrast to the positive correlation observed when using positive predictions ($\mathsf{Conf}^{+}$). Although the proposed fixed-threshold approach (0.3) occasionally outperforms the adaptive method, the performance gap remains marginal. Moreover, the fixed threshold consistently exhibits lower correlation with UP-DETR, for which the optimal threshold has been empirically determined to be approximately 0.5.
@@
$\mathsf{ContrastiveConf}$ (Optimal) 0.700 0.684 0.709 0.695 0.601 0.633 0.662 0.634 0.612 0.657 0.672 0.643
@@
Ours $\mathsf{Conf}^{+}$ (OCE) 0.442 0.477 0.547 0.512 0.237 0.355 0.393 0.370 0.228 0.291 0.270 0.274
$\mathsf{Conf}^{-}$ (OCE) -0.633 -0.581 -0.571 -0.585 -0.365 -0.389 -0.379 -0.394 -0.245 -0.325 -0.253 -0.283
$\mathsf{ContrastiveConf}$ (OCE) 0.585 0.570 0.597 0.587 0.334 0.432 0.431 0.438 0.287 0.362 0.302 0.327
$\mathsf{ContrastiveConf}$ (0.3) 0.568 0.574 0.599 0.594 0.297 0.437 0.430 0.464 0.248 0.346 0.296 0.325
$\mathsf{ContrastiveConf}$ (Top-100) -0.619 -0.253 -0.490 -0.321 -0.409 -0.342 -0.358 -0.360 -0.262 -0.175 -0.215 -0.139
(a) AP
Oracles $\mathsf{Conf}^{+}$ (Optimal) 0.488 0.591 0.637 0.611 0.584 0.638 0.616 0.642 0.598 0.640 0.624 0.639
$\mathsf{Conf}^{-}$ (Optimal) -0.583 -0.580 -0.555 -0.575 -0.248 -0.305 -0.316 -0.324 -0.137 -0.185 -0.175 -0.156
$\mathsf{ContrastiveConf}$ (Optimal) 0.675 0.664 0.677 0.673 0.594 0.630 0.629 0.646 0.629 0.634 0.629 0.632
Baselines $\mathsf{Conf}^{+}$ (Top-100) -0.586 -0.581 -0.546 -0.571 -0.375 -0.369 -0.351 -0.394 -0.216 -0.217 -0.196 -0.204
$\mathsf{Conf}^{+}$ (0.3) 0.445 0.435 0.494 0.465 0.229 0.321 0.342 0.379 0.203 0.223 0.179 0.194
Ours $\mathsf{Conf}^{+}$ (OCE) 0.426 0.443 0.501 0.471 0.242 0.331 0.349 0.364 0.256 0.239 0.198 0.211
$\mathsf{Conf}^{-}$ (OCE) -0.600 -0.576 -0.554 -0.573 -0.317 -0.394 -0.357 -0.406 -0.194 -0.277 -0.216 -0.247
$\mathsf{ContrastiveConf}$ (OCE) 0.560 0.545 0.554 0.552 0.314 0.416 0.388 0.438 0.285 0.302 0.229 0.263
$\mathsf{ContrastiveConf}$ (0.3) 0.532 0.553 0.557 0.561 0.277 0.418 0.388 0.456 0.237 0.298 0.216 0.259
$\mathsf{ContrastiveConf}$ (Top-100) -0.586 -0.189 -0.435 -0.250 -0.375 -0.315 -0.332 -0.340 -0.216 -0.122 -0.171 -0.121
(b) AP@50
Oracles $\mathsf{Conf}^{+}$ (Optimal) 0.490 0.593 0.629 0.598 0.486 0.586 0.608 0.563 0.488 0.607 0.635 0.574
$\mathsf{Conf}^{-}$ (Optimal) -0.572 -0.541 -0.521 -0.538 -0.280 -0.244 -0.284 -0.274 -0.154 -0.234 -0.182 -0.189
$\mathsf{ContrastiveConf}$ (Optimal) 0.672 0.649 0.662 0.649 0.541 0.562 0.614 0.563 0.534 0.627 0.641 0.587
Baselines $\mathsf{Conf}^{+}$ (Top-100) -0.580 -0.555 -0.523 -0.549 -0.373 -0.314 -0.325 -0.348 -0.226 -0.266 -0.206 -0.216
$\mathsf{Conf}^{+}$ (0.3) 0.459 0.434 0.497 0.467 0.214 0.320 0.348 0.351 0.188 0.231 0.290 0.241
Ours $\mathsf{Conf}^{+}$ (OCE) 0.423 0.447 0.503 0.478 0.234 0.330 0.361 0.322 0.206 0.272 0.293 0.262
$\mathsf{Conf}^{-}$ (OCE) -0.597 -0.539 -0.520 -0.537 -0.339 -0.325 -0.327 -0.342 -0.220 -0.321 -0.217 -0.269
$\mathsf{ContrastiveConf}$ (OCE) 0.557 0.532 0.548 0.545 0.319 0.386 0.390 0.381 0.258 0.346 0.314 0.312
$\mathsf{ContrastiveConf}$ (0.3) 0.543 0.535 0.551 0.548 0.279 0.384 0.385 0.406 0.233 0.324 0.316 0.306
$\mathsf{ContrastiveConf}$ (Top-100) -0.580 -0.225 -0.435 -0.291 -0.373 -0.295 -0.320 -0.334 -0.226 -0.170 -0.188 -0.137
(c) AP@75
@@
Table 8: Comparison of post-processing schemes and baseline methods based on their Pearson correlation coefficients with image-level reliability when $\mathsf{Conf}^{+}$ and $\mathsf{ContrastiveConf}$ are applied, respectively. We evaluated three standard methods: confidence thresholding, top-$k$, and NMS (without confidence thresholding). For each scheme, the optimal hyperparameter selected from the validation set is shown in parentheses, and is applied on the test set. the strongest correlations are highlighted in bold. Confidence thresholding achieves the lowest OCE, demonstrating its better efficacy compared to the other schemes. For Deformable-DETR, Cal-DETR, and DINO, the optimal thresholds are approximately $0.3$, aligning well with values commonly employed in previous studies. For UP-DETR, the optimal thresholds exceed $0.5$, highlighting the potential limitation of using a fixed threshold in image-level uncertainty quantification (UQ) applications.
@@
NMS -0.588 (0.10) -0.560 (0.10) -0.548 (0.10) -0.562 (0.10) -0.402 (0.10) -0.346 (0.10) -0.332 (0.10) -0.365 (0.10) -0.240 (0.10) -0.265 (0.10) -0.218 (0.10) -0.224 (0.10)
$\mathsf{ContrastiveConf}$ Thresholding 0.597 (0.55) 0.574 (0.30) 0.599 (0.30) 0.594 (0.30) 0.408 (0.80) 0.437 (0.30) 0.443 (0.20) 0.472 (0.35) 0.312 (0.75) 0.362 (0.25) 0.315 (0.40) 0.352 (0.35)
Top-$k$ 0.629 (1.00) 0.506 (1.00) 0.474 (1.00) 0.527 (1.00) 0.409 (1.00) 0.377 (1.00) 0.324 (1.00) 0.411 (1.00) 0.284 (1.00) 0.254 (1.00) 0.156 (1.00) 0.298 (1.00)
NMS -0.408 (0.10) 0.585 (0.10) 0.535 (0.10) 0.574 (0.10) -0.101 (0.10) 0.368 (0.10) 0.347 (0.10) 0.373 (0.10) -0.001 (0.10) 0.262 (0.10) 0.228 (0.10) 0.239 (0.10)
(a) Predictions for a High-Reliability Prediction: Horse
(b) Predictions for a High-Reliability Prediction: Chair
(c) Predictions for a Low-Reliability Prediction: Chair
Figure 5: The evolution of predictions made by Cal-DETR across decoder layers for a high-reliability and low-reliability image. The positive query (indexed by 0 and bordered with green) and the five negative queries with the largest IoU (indexed by 1-5 and bordered with red) are presented.
(a) Predictions for a Low-Reliability Object: Laptop
(b) Predictions for a High-Reliability Object: Dog
Figure 6: The evolution of predictions made by Cal-DETR across decoder layers for two different objects (one reliable and one unreliable) in a low-reliability image. The positive query (indexed by 0 and bordered with green) and the five negative queries with the largest IoU (indexed by 1-5 and bordered with red) are presented.
(a) UP-DETR (COCO)
(b) D-DETR (COCO)
(c) Cal-DETR (COCO)
(d) DINO (COCO)
(e) UP-DETR (Cityscapes)
(f) D-DETR (Cityscapes)
(g) Cal-DETR (Cityscapes)
(h) DINO (Cityscapes)
(i) UP-DETR (Foggy)
(j) D-DETR (Foggy)
(k) Cal-DETR (Foggy)
(l) DINO (Foggy)
Figure 7: Impact of confidence threshold selection on various performance metrics in UP-DETR, Deformable-DETR, Cal-DETR, and DINO on COCO, Cityscapes, and Foggy Cityscapes.
(a) COCO
(b) Cityscapes
(c) Foggy Cityscapes
Figure 8: Impact of the scaling factor ($\lambda$) on image-level UQ performance of $\mathsf{ContrastiveConf}$. Pearson correlation coefficient (PCC) using various scaling factors is reported. The optimal scaling factor lies within the range of $5.0$ to $10.0$, while this range generalizes well across out-of-distribution datasets. Furthermore, it shows the efficacy of $\mathsf{ContrastiveConf}$ over $\mathsf{Conf}^{+}$ (i.e., $\mathsf{ContrastiveConf}$ with $\lambda=0.0$).

### v2

@@ V-C Numerical Analysis: Effectiveness of OCE
To showcase OCE’s effectiveness in assessing calibration quality, we analyze how different metrics rank calibration errors across four DETR models. As baselines, we measure ECEs using the top $100$ detections  a standard confidence threshold of 0.3, and the threshold that minimizes the different metrics on the validation set. In a similar manner, we assess our OCEs using thresholds that minimize each OCE metric on the validation set. For a fair comparison, we execute the optimal matching process on each test dataset to obtain the optimal positive set. We then evaluate D-ECE, LA-ECE0, and OCE using these sets as reference scores and calculate the PCC correlation to the other scoring schemes across different datasets.
Table 1 presents the correlations between the three calibration metrics on the optimal positive set and various methods. The first three rows demonstrate that the calibration metrics—D-ECE, LA-ECE, and our proposed OCE—are highly correlated. This indicates that all three metrics effectively capture the notion of calibration quality. Additionally, these metrics show a decent correlation with AP metrics. It is important to note that AP primarily accounts for accuracy rather than calibration quality, which may explain why the correlation, particularly for AP@50 on the out-of-distribution dataset, is not perfectly aligned.
LRP, D-ECE, and LA-ECE0 exhibit inverse correlations with calibration quality when used alone. Furthermore, LA-ECE0 with a non-optimal positive set also shows an inverse correlation with the overall calibration errors estimated on the optimal positive set. This supports our assertion that these metrics are inadequate for measuring models’ calibration quality in conjunction with the employed post-processing scheme, unlike our OCE.
Another important attribute of OCE is that it can not only compare different models but also identify the most reliable configuration of the same model. As illustrated in Figure 5, the confidence threshold vs. OCE curve exhibits a bell shape, unlike ECE metrics, and the optimal OCE is achieved around a threshold of $0.3$. This is particularly notable because it reflects the practical choices in many calibration studies to date [12, 20], affirming the reliability of using OCE.
The D-ECE and OCE scores measured at the optimal LRP threshold—denoted by D-ECE (LRP) and OCE (LRP), respectively—show a strong correlation to those computed from the optimal positive set; thus, the method of using LRP to identify a positive set [20] appears acceptable. Experimentally, we confirmed that OCE and LRP provide thresholds in a similar range across different settings.
Nonetheless, the biggest advantage of using OCE over other metrics is that OCE measures the model’s calibration quality alongside the post-processing scheme employed. While measuring the calibration error on the optimal positive set would ideally assess quality, this optimal set is not accessible during test-time inference. Consequently, the score measured on the optimal set may not correspond to the quality experienced in real-world applications, such as those employing confidence thresholding schemes.
Using a fixed separation threshold of $0.3$ might serve as a reasonable approximation for distinguishing between optimal positives and negatives. However, trained or post-calibrated DETR models often have varying optimal confidence thresholds. For example, UP-DETR has an optimal threshold of approximately $0.5$, and low-temperature calibrated models are likely to require even higher thresholds. Consequently, employing a fixed threshold poses potential risks. Similarly, using a fixed number of top confident samples (e.g., Top-$100$) is inadequate, as it may include more predictions than the actual number of ground truth positives.
VI Quantifying Image-level Reliability
As demonstrated, positive and negative predictions exhibit varying levels of reliability. Interestingly, having predictions with low confidence scores does not necessarily imply low reliability. Our empirical observations show that confidence scores in negative predictions are actually inversely correlated with image-level reliability. More specifically, for a reliable instance, we observe that the confidence of positive predictions increases across the decoder layers, while that of negative predictions remains low; this results in a large gap between positive and negative predictions. In contrast, for unreliable instances (e.g., Figure 3a), the confidence score of the positive prediction does not increase across the layers, whereas the scores of negative predictions are either slightly elevated or remain unchanged. Consequently, there is a small gap between positive and negative predictions (e.g., Figure 3b).
VI-A Proposed Method: Quantifying Reliability by Contrasting
$\mathsf{ContrastiveConf}(\bm{x})=\mathsf{Conf}^{+}(\bm{x})-\mathsf{Conf}^{-}(\bm{x})$ (11)
$\mathsf{Conf}^{+}(\bm{x})=\frac{1}{|\hat{\mathcal{D}}_{\theta}^{+}(\bm{x})|}\sum\nolimits_{(\hat{\bm{p}},\hat{\bm{b}})\in\hat{\mathcal{D}}_{\theta}^{+}(\bm{x})}\max_{c}\hat{\bm{p}}[c]$ (12)
$\mathsf{Conf}^{-}(\bm{x})=\frac{1}{|\hat{\mathcal{D}}_{\theta}^{-}(\bm{x})|}\sum\nolimits_{(\hat{\bm{p}},\hat{\bm{b}})\in\hat{\mathcal{D}}_{\theta}^{-}(\bm{x})}\max_{c}\hat{\bm{p}}[c]$ (13)
where $\hat{\mathcal{D}}_{\theta}^{+}(\bm{x})$ and $\hat{\mathcal{D}}_{\theta}^{-}(\bm{x})$ are predicted sets of positive and negative predictions, respectively, and $\lambda$ is a scaling factor that can be determined on the validation set. We include an ablation study of the scaling factor in the subsequent section. Our results demonstrate that the proposed method is robust to the choice of scaling factor and consistently outperforms baseline approaches, achieving the robust performance with a scaling factor of $5.0$ $10.0$.
Identifying a $\hat{\mathcal{D}}_{\theta}^{+}(\bm{x})$ from $\hat{\mathcal{D}}_{\theta}(\bm{x})$ is a crucial factor in the success of this approach. In practice, however, neither the ground truth nor the optimal positives are available during the test time. Instead, we approximate the ground truth separation by applying a post-processing algorithm $\mathcal{A}^{*}$ that minimizes the calibration error on the validation set:
$\hat{\mathcal{D}}_{\theta}^{+}(\bm{x})=\mathcal{A}^{*}\circ\hat{\mathcal{D}}_{\theta}(\bm{x})$ (14)
$\mathcal{A}^{*}=\text{argmin}_{\mathcal{A}}\mathsf{OCE}(\mathcal{A}\circ\hat{\mathcal{D}}_{\theta};\mathcal{I}_{val})$ (15)
where $\mathcal{I}_{val}=\cup_{i\in\mathcal{V}}\{(i,j)\}_{j=1}^{N_{i}}$ is the set of all objects indices in the validation dataset $\mathcal{V}$.
@@ VI-A Proposed Method: Quantifying Reliability by Contrasting
Table 2: Comparison of the proposed method ($\mathsf{ContrastiveConf}$) with baseline methods on their Pearson correlation coefficient with the image-level reliability in Equation (5), with different models. $\mathsf{Perf}$ is evaluated using AP. For each model, we apply the optimal matching, the standard post-processing scheme (Top-$100$ and confidence thresholding by $0.3$), or the proposed OCE-based post-processing scheme, as indicated within the parentheses. The strongest correlations are highlighted in bold, while negative correlations appear in red. First, the proposed contrasting method, $\mathsf{ContrastiveConf}$, achieves the highest correlation with image-level reliability. In addition, it is noteworthy that the average confidence of negative predictions ($\mathsf{Conf}^{-}$) exhibits a negative correlation with image-level reliability, in contrast to the positive correlation observed when using positive predictions ($\mathsf{Conf}^{+}$).
@@
$\mathsf{ContrastiveConf}$ (Optimal) 0.700 0.648 0.684 0.662 0.601 0.659 0.656 0.644 0.612 0.660 0.667 0.647
@@
Ours $\mathsf{Conf}^{+}$ (OCE) 0.477 0.478 0.547 0.512 0.252 0.330 0.393 0.370 0.191 0.276 0.270 0.274
$\mathsf{Conf}^{-}$ (OCE) -0.640 -0.575 -0.571 -0.585 -0.349 -0.382 -0.379 -0.394 -0.256 -0.322 -0.253 -0.283
$\mathsf{ContrastiveConf}$ (OCE) 0.656 0.588 0.628 0.613 0.359 0.407 0.441 0.440 0.275 0.350 0.317 0.327
VI-B Numerical Analysis
We compare different methods for quantifying per-image reliability based on different separation methods. For separation methods, we apply the optimal matching (as a reference) as well as the practical post-processing schemes such as fixed and adaptive confidence thresholding. We measure the Pearson correlation coefficient of each method with the $\mathsf{ImReli}$ computed based on AP metric (see Appendix for the detail).
Table 2 shows that the proposed $\mathsf{ContrastiveConf}$ consistently achieves the best correlation with image-level reliability. Interestingly, the absolute value of $\mathsf{Conf}^{-}$ often surpasses that of $\mathsf{Conf}^{+}$ when a non-optimal separation scheme is applied. This observation highlights how our contrasting approach, which leverages the strengths of both $\mathsf{Conf}^{+}$ and $\mathsf{Conf}^{-}$, achieves robust performance across diverse settings.
(a) COCO
(b) Cityscapes
(c) Foggy Cityscapes
Figure 6: Impact of the scaling factor ($\lambda$) on image-level UQ performance of $\mathsf{ContrastiveConf}$ (OCE). Pearson correlation coefficient (PCC) using various scaling factors is reported. The optimal scaling factor lies within the range of $5.0$ to $10.0$, while this range generalizes well across out-of-distribution datasets. Furthermore, it shows the efficacy of $\mathsf{ContrastiveConf}$ over $\mathsf{Conf}^{+}$ (i.e., $\mathsf{ContrastiveConf}$ with $\lambda=0.0$).
In addition, we conduct an ablation study on the scaling factor ($\lambda$), with results presented in Figure 6. The study reveals that the best performing scaling factor for COCO dataset lies between $5.0$ and $10.0$. Notably, this range remains effective even for out-of-distribution datasets such as Cityscapes and Foggy Cityscapes. Furthermore, the results demonstrate that $\mathsf{ContrastiveConf}$ with $1.0\leq\lambda\leq 10.0$ consistently outperforms $\mathsf{Conf}^{+}$ (i.e., $\lambda=0.0$, the leftmost point on each line). However, we also observe a rapid drop in performance when the scaling factor is excessively large, particularly on the OOD dataset, Foggy Cityscapes. Thus, caution is needed when applying to substantially different datasets.
@@ VI-B Numerical Analysis
Table 4: Comparison of post-processing schemes and baseline methods based on their Pearson correlation coefficients with image-level reliability when $\mathsf{Conf}^{+}$ and $\mathsf{ContrastiveConf}(\lambda=1.0)$ are applied, respectively. We evaluated three standard methods: confidence thresholding, top-$k$, and NMS (without confidence thresholding). For each scheme, the optimal hyperparameter selected from the validation set is shown in parentheses, and is applied on the test set. the strongest correlations are highlighted in bold. Confidence thresholding achieves the lowest OCE, demonstrating its better efficacy compared to the other schemes. For Deformable-DETR, Cal-DETR, and DINO, the optimal thresholds are approximately $0.3$, aligning well with values commonly employed in previous studies. For UP-DETR, the optimal thresholds often exceed $0.5$, highlighting the potential limitation of using a fixed threshold for image-level UQ.
@@
NMS -0.524 (0.10) -0.508 (0.10) -0.488 (0.10) -0.510 (0.10) -0.370 (0.10) -0.324 (0.10) -0.273 (0.10) -0.338 (0.10) -0.205 (0.10) -0.251 (0.10) -0.202 (0.15) -0.204 (0.10)
$\mathsf{ContrastiveConf}$ Thresholding 0.597 (0.55) 0.513 (0.25) 0.563 (0.25) 0.540 (0.25) 0.408 (0.80) 0.389 (0.30) 0.422 (0.20) 0.437 (0.35) 0.312 (0.75) 0.317 (0.25) 0.288 (0.15) 0.311 (0.35)
Top-$k$ 0.629 (1.00) 0.268 (1.00) 0.320 (1.00) 0.290 (1.00) 0.409 (1.00) 0.169 (1.00) 0.151 (1.00) 0.220 (1.00) 0.284 (1.00) 0.077 (1.00) 0.060 (1.00) 0.174 (1.00)
NMS -0.227 (0.10) -0.157 (0.90) -0.402 (0.10) -0.140 (0.90) -0.044 (0.80) -0.034 (0.75) -0.203 (0.10) -0.080 (0.80) 0.035 (0.75) -0.035 (0.75) -0.149 (0.90) -0.075 (0.70)
(a) Thresholding
(b) Top-$k$
(c) NMS
Figure 7: Impact of parameter selection on OCE (y-axis inverted) and the Pearson correlation coefficient (PCC) between $\mathsf{Conf}^{+}$ and image-level reliability in Cal-DETR on COCO for different post-processing schemes.
VII Comparative Study on Post-Processing
In this section, we provide a comprehensive analysis regarding the impact of post-processing on the model’s overall calibration quality and UQ performance. For the post-processing (i.e., separation) methods, we compare the following approaches: (1) applying a threshold on the confidence score, (2) selecting the Top-$k$ predictions, and (3) utilizing NMS.
Primarily, we evaluate OCE by varying the hyperparameter for each method (i.e., confidence threshold, $k$, and IoU threshold, respectively) and compare the resulting performance. In this first analysis, to exclude the dependency of the final performance on OCE, we choose the best hyperparameter on the validation set for each setting and compare their best possible performances. Table 3 shows that the confidence thresholding approach outperforms the top-$k$ and NMS approaches. The top-$k$ approach underperforms because the number of objects in an image varies significantly, making it impractical to determine a single optimal value for $k$. Top-$k$ becomes optimal when $k=20$. Given that the average and 95th percentile number of objects per image in the COCO dataset are 7 and 22, respectively, these results appear reasonable. Therefore, we confirm that using an excessively large number (e.g., 100) for top-$k$ is inadequate for achieving well-calibrated predictions. On the other hand, with NMS (when applied without preceding confidence thresholding), we empirically observe that although most of the optimal positive queries are included, a substantial number of negative queries are retained, often outnumbering the positive ones by several times.
Similarly, we compare the image-level uncertainty quantification performance, and the results are shown in Table 4. In this experiment, we fixed the scaling factor $\lambda$ at $1.0$ to minimize its influence on selecting the optimal hyperparameter. We reconfirm the effectiveness of confidence thresholding; for Deformable-DETR, Cal-DETR, and DINO, the optimal thresholds are approximately $0.3$, aligning well with values commonly employed in previous studies. For UP-DETR, the optimal thresholds exceed $0.5$, highlighting the potential limitation of using a fixed threshold in image-level UQ applications. Moreover, the correlation is often negative when NMS is applied along with the $\mathsf{Conf}^{+}$ framework. This is because, while most of the optimal positive queries are likely to be included with NMS, a substantial number of optimal negative queries remain within the final subset, often outnumbering the positive ones by several times. As a result, applying NMS leads to an inaccurate reliability assessment. However, many schemes, including NMS, achieve significant improvement when used with the proposed contrastive framework; even if the post-processing scheme is inaccurate, the framework can robustly perform by leveraging the negative predictions.
Lastly, we plot the OCE and image-level UQ performance with different hyperparameter selections to show the sensitivity of each scheme to the choice of hyperparameter. Figure 7 highlight that carefully identifying a reliable subset is crucial for achieving both high object-level calibration quality and effective image-level uncertainty quantification performance. We also present this with exemplary visualizations in Figure 8 and in the Appendix.